OpenStudy (anonymous):

how to solve -x+2/7=2x+86/7

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

-3x=84/7 -3x=12 x=-4

OpenStudy (anonymous):

-3x=84/7 -3x=12 x=-4

OpenStudy (anonymous):

This is setting two equations equal to each other to find the point where their lines intersect each other. They're both in y=mx+b form, so you just have to solve the equation algebraically -X+2/7=2X+86/7 -X+X+2/7=2X+X+86/7 2/7=3X+86/7 2/7-86/7=3X+86/7-86/7 -84/7=3X -12=3X -12=3X --- --- 3 3 -4=X When you have your X-value for the point, you can plug it into either of the two equations you started with and solve for Y: y=-x+2/7 y=2x+86/7 Now you have your coordinate!